OpenStudy (anonymous):

1. TRAVEL During Sonia’s trip across the country, she traveled 2,884 miles. Her trip took 7 days. Find a unit rate to represent the average miles she traveled per day during the trip. 2. BUDGET Steve was offered $5,025 per year for a weekend lifeguarding job at a local pool. He wants to know how much his monthly income will be at this salary level. What is his rate of pay in dollars per month? 3. MUSIC Randall recorded 8 songs on his most recent CD. The total length of the CD is 49 minutes. Find a unit rate to represent the average length per song on the CD. 4. CARPETING Hana paid $1,200 for

OpenStudy (anonymous):

1) \[Vav = \Delta D/ T\] D= distance T= days Average Distance travelled = D/T 2) Salary = 5,025 Months in a year = 12 Total pay period per month = Salary/months 3) Cd lenth = 49 Tracks 8 Average lenth = Cd/Track